OpenStudy (anonymous):

Evaluate the following expression. 2^-3

OpenStudy (anonymous):

A negative exponent means that it is on the bottom of the fraction. Since this doesn't have a coefficient, it's understood that the coefficient is 1. In an early problem, we saw that an exponent of -1 meant 1/x. So what would an exponent of -3 translate to?

OpenStudy (anonymous):

2^-3=1/2^3=1/8=0.125

OpenStudy (anonymous):

well you have to evaluate it as a fraction..

OpenStudy (anonymous):

i did not get i solve it in fraction form 1/8 and in decimal t00 0.125

OpenStudy (anonymous):

Doing it on a TI-83 plus calculator gives .125